OpenStudy (anonymous):

Find the quotient, 2^3 ÷ 5^2 10^1?.

OpenStudy (anonymous):

No...

OpenStudy (anonymous):

do u knw 2^3=?

OpenStudy (theviper):

8/25

OpenStudy (anonymous):

2^3 = 2*2*2=8

OpenStudy (anonymous):

5^2 = 5*5=25

OpenStudy (anonymous):

oh oops

OpenStudy (anonymous):

Similarly 5^2 = 5*5 = 25..

OpenStudy (anonymous):

so 8/25

OpenStudy (anonymous):

u got it

OpenStudy (anonymous):

Yes..